# SAFE: Multitask Failure Detection for Vision-Language-Action Models

### v1

@@ 7 Conclusion
Limitations: Most recent VLAs have shown capabilities in handling diverse modalities, controlling diverse embodiments and learning latent actions from non-robotic action-less video data [66, 67]. This paper only considers multitask failure detection for manipulation tasks, and it’s not clear how well the failure detectors generalize across embodiments, sim2real or to the action-less videos. Besides, SAFE only uses features from the last layer and how to effectively aggregate information across multiple layers of a VLA remains future work.
@@
The authors were partially supported by the Toyota Research Institute (TRI) and NSERC Discovery Grant. The authors thank Blerim Abdullai, Sebastian Aegidius, Jasper Gerigk, Ruthrash Hari, and Wei-Cheng Tseng for helpful discussions and feedback.
@@
Appendix A Experiment Details
A.1 Vision-Language-Action Models
@@
For all VLA models, we ablate different methods of aggregating the hidden features $E$ into a single feature vector $\mathbf{e}$ and select the best method according to $\mathcal{D}_{\text{eval-seen}}$ performance. The detailed ablation results are shown in Section A.7.
@@
A.2 SAFE Failure Detector
@@
A.3 Failure Detection Baselines
@@ A.5 Benchmark Details
Note that the LIBERO simulator stops the rollout execution when the robot finishes the task (considered a success) or a maximum rollout length is reached (considered a failure). Therefore, in the generated rollouts, failed ones always have the maximum length, but successful ones are shorter. This could result in an unfair advantage for some of the compared failure detectors (if a failure detector simply learns to count the time elapsed, i.e., $s_{t}=t$, it will achieve perfect failure detection since failed rollouts have a fixed and longer duration). To ensure a fair comparison, for evaluation in Table 1, we compute the minimum rollout length for each task and use that as $T$ for that task. The failure detection performance (in ROC-AUC) is then determined based on $s_{T}$, where $T$ is the same for all successful and failed rollouts within each task.
@@
On this benchmark, we adopt the pretrained model checkpoints of $\pi_{0}^{*}$ [62]. Note that $\pi_{0}^{*}$ model checkpoints are trained separately on the Google Robot embodiment and the WidowX embodiment, which results in two model checkpoints that have different internal feature spaces. Therefore, all failure detectors are trained and evaluated on each embodiment separately as well. All reported evaluation metrics are computed separately for each embodiment and then averaged. In Table 2, we list the tasks used for failure detection on SimplerEnv. We generate 100 rollouts for each task with random initial configurations, and the success rates of $\pi_{0}^{*}$ on each task are also listed in Table 2. A rollout stops after the maximum number of allowed timesteps have passed, regardless of task success or failure. Within each embodiment, 1 out of 4 tasks is unseen, and within the seen tasks, 66% of the rollouts are in $\mathcal{D}_{\text{train}}$ and the remaining 33% in $\mathcal{D}_{\text{eval-seen}}$.
@@
Table 2: List of tasks used in SimplerEnv benchmark.
Table 3: List of tasks used in the real-world experiments.
@@
Real-world experiments with Franka robot: In Table 3, we list the tasks used in the real-world experiments. For each task, we set a number of timesteps $T$ allowed for one rollout, and all rollouts of the same task are terminated after the same $T$ timesteps regardless of task success or failure. In Fig. 6, we further visualize some example successful and failed rollouts from the real-world experiments.
@@
A.6 Training Details
@@ A.7 Hyperparameter Tuning
To determine the hyperparameters for the proposed SAFE and baselines, we perform a grid search over them and select the ones with the highest failure detection performance (ROC-AUC) on the $\mathcal{D}_{\text{eval-seen}}$ split. In Table 5, Table 6, and Table 7, we report the hyperparameters we have searched over and the values with the best performance. Note that for the real-world experiments, we fix the $\mathbf{e}_{t}$ to be the “pre-logits” with “Mean” aggregation.
Appendix B Additional Results and Discussions
B.1 Feature Visualization and Analysis
@@
We perform the feature analysis similar to Section 4.1 and Fig. 1 on other benchmarks and show the plots in Fig. 7, Note that in this feature analysis process, the t-SNE algorithm was performed on the VLA’s embeddings without any learning. Therefore, the feature dimension reduction process is unsupervised and does not know about task successes or failures.
@@
Comparing the plots in Fig. 1 and Fig. 7, we can see that the embedding spaces from VLAs are different from each other, which corresponds to the different failure patterns presented by the VLAs. For $\pi_{0}$-FAST (Fig. 1) and $\pi_{0}$ on LIBERO (Fig. 7a and b), when task execution fails, the embeddings fall into the same region (“failure zone”). This corresponds to the major failure mode of $\pi_{0}$-FAST trained on the LIBERO dataset, where the predicted actions $\mathbf{A}_{t}$ become unstable and the robot arms move to weird configurations and out of the observation frame. For OpenVLA on LIBERO (Fig. 7c and d). we observe that for most failed rollouts, the robot freezes at or shakes around certain configurations during the middle of task execution. Such failed rollouts result in features very close to each other, which corresponds to small blobs of red dots in Fig. 7c.
@@
Despite the different appearances of the embedding spaces from the above benchmarks, their successful and failed rollouts are separable in the feature space. This is aligned with the high performance of the proposed SAFE and the embedding-based baseline methods. Moreover, although the embeddings of the failed rollouts from OpenVLA are spread over the space and do not form a unified “failure zone”, SAFE is still able to learn to separate task failures from successes (possibly by extracting the correlations that are not visualized by t-SNE) and generalize well to unseen tasks, as reported in Table 1.
@@
However, the visualized embeddings of $\pi_{0}$-FAST on the real Franka robot (Fig. 7e and f) are different, where embeddings from successful and failed rollouts are not easily separable through the t-SNE visualization. We hypothesize that because the tasks we used for real-world experiments are more diverse, their failures do not have a unified semantic meaning, and thus the embeddings are not clearly separated in the visualization. This explains the limited performance of all failure detection methods as reported by Fig. 5, where ROC-AUC is at most 64 on $\mathcal{D}_{\text{eval-unseen}}$. Nevertheless, SAFE-MLP still outperforms all baselines on both seen and unseen splits in this evaluation.
@@
B.2 Conformal Prediction Results
@@
We use functional CP [8, 22] to determine the time-varying thresholds $\delta_{t}$ for failure detection. By varying the significance level $\alpha$ used in functional CP, we can adjust the conservativeness of failure detection and get different performance. In Fig. 8, we plot the change of TNR (True Negative Rate), TPR (True Positive Rate) and Bal-acc (Balanced Accuracy, $\frac{\text{TNR}+\text{TPR}}{2}$) w.r.t. $\alpha$.
@@
Note that we calibrate the CP bands on successful rollouts (negative data points), and thus if the assumptions used in CP ($s_{t}$ are sampled i.i.d.) hold, the TNR rate is lower bounded by and close to $1-\alpha$ (the gray dashed line in the TNR plots in Fig. 8[22]. However, as the multitask failure detection problem requires detecting failures on tasks that are not in the training or the calibration sets, we need to calibrate CP bands on $\mathcal{D}_{\text{eval-seen}}$ and then evaluate them on $\mathcal{D}_{\text{eval-unseen}}$. Therefore, the i.i.d assumption may not hold, and TNR may deviate from the gray dashed line.
@@
From Fig. 8, we can see that on OpenVLA+LIBERO and $\pi_{)}^{*}$+SimplerEnv benchmarks, the TNR curves obtained by SAFE are close to the gray dashed line $1-\alpha$, while those on the other 3 benchmarks are lower than $1-\alpha$. A similar phenomenon is also observed for the baseline methods: none of the TNR curves obtained from the baselines consistently conform to the $1-\alpha$ curve across all benchmarks. We attribute this to the challenging nature of the multitask failure detection problem, where the failure scores for calibration and evaluation may not come from the same distribution. Nevertheless, we still adopt the functional CP as a principled method to determine the time-varying failure detection threshold $\delta_{t}$. Moreover, from Fig. 8, we can see that SAFE can achieve higher TPR and result in fewer false negatives compared to the baselines. This is crucial for safety-critical environments, where a missing failure (false negative) can be much more catastrophic than a false alarm (false positive).
@@
Figure 8: Additional failure detection results using $\delta_{t}$ obtained by functional CP. These plots show TNR (left column), TPR (middle column), and Bal-acc (right column) w.r.t. the significance level $\alpha$, for each evaluation benchmark. These plots are obtained with random seed$=0$.
B.3 Failure Detection Time
@@
As mentioned in Section 6.2, we manually label when the failure happens for the failed rollouts. The labeling process is based on video recordings after all rollouts are collected and no interventions were done during the task execution. While the exact times of failure are clear for some failure modes (e.g. dangerous actions, breaking objects), they can be ambiguous and hard to annotate for other failure modes. For example, a policy may freeze in the middle of task execution, and after that either recovering from it or getting stuck indefinitely can be possible. In another case, a policy may repeatedly try grasping the object but keep missing the grasp until timeout, and it’s hard to determine a single point of failure. To handle such cases, we instruct the human annotators to pick the time where they think intervention is needed and they should take over control to prevent an execution failure. In practice, for the above ambiguous failure modes, we annotate the failures after the policy gets stuck by a few seconds or re-tries the grasping action a few times. For some rollouts that look very plausible but do not succeed due to the time limit, the failure time is annotated as the end of the rollout. Note that we annotate only the failed rollouts and not the successful ones, even though they may also show subtle signs of failure in the middle.
@@
In Fig. 9, we compare the times of failure detected by the proposed SAFE-MLP model and a human annotator. From Fig. 9, we can see that for both $\pi_{0}$ and $\pi_{0}$-FAST models, SAFE-MLP can detect failures before they happen (as identified by a human). When used for $\pi_{0}$-FAST deployed on LIBERO, SAFE-MLP can forecast failures well in advance and even predict 40% of the failures after the first timestep.
@@
Furthermore, from Fig. 9a and Fig. 9c, we can see that the blue curves jump up on the right edge of the plots. This means that the human annotator does not think these rollouts are failures until the very last moment, where the VLA model is probably on the right track and fails only due to timeout. We think such failures are also hard for failure detectors to detect, and it explains the low performance of all failure detectors on these benchmarks.
@@ B.4 Result Variance
In Table 4, we report the standard deviation for all results in Table 1 and Fig. 5 left. Note that for the repeated runs, not only are they using different random seeds, also the tasks are split differently into the seen and the unseen subsets. Since different tasks have different difficulties for failure detection, it is normal to see large standard deviations in Table 4. From Table 4, we can see that the proposed SAFE methods achieve high averaged performance with relatively low standard deviations compared to the baselines, across all evaluation benchmarks. This signifies the strong and also stable performance of SAFE.

### v2

@@ 7 Conclusion
Limitations: Most recent VLAs have shown capabilities in handling diverse modalities, controlling diverse embodiments, and learning latent actions from non-robotic action-less video data [69, 70]. This paper only considers multitask failure detection for manipulation tasks, and it is not clear how well the failure detectors generalize across embodiments, sim2real or to action-less videos. Besides, SAFE only uses features from the last layer, and how to effectively aggregate information across multiple layers of a VLA remains an open question for future work.
@@
The authors were partially supported by the Toyota Research Institute (TRI) and NSERC Discovery Grant. The authors thank Blerim Abdullai, Sebastian Aegidius, Jasper Gerigk, Ruthrash Hari, and Wei-Cheng Tseng for helpful discussions and feedback. The authors also thank the anonymous reviewers and area chairs for their constructive feedback.
@@
Appendix A Potential Societal Impact
This work advances the safety and reliability of VLAs through multitask failure detection, which can reduce unintended behaviors during robot deployment. However, the proposed framework could be misused in surveillance or fully autonomous systems with limited human oversight. Additionally, biases or privacy issues may arise from training data collected during robot interactions. These risks can be mitigated through responsible data handling, transparency in model release, and maintaining human oversight in downstream applications.
Appendix B Experiment Details
B.1 Vision-Language-Action Models
@@
For all VLA models, we ablate different methods of aggregating the hidden features $E$ into a single feature vector $\mathbf{e}$ and select the best method according to $\mathcal{D}_{\text{eval-seen}}$ performance. The detailed ablation results are shown in Section˜B.8.
@@
B.2 SAFE Failure Detector
@@
B.3 Failure Detection Baselines
@@ B.5 Benchmark Details
Note that the LIBERO simulator stops the rollout execution when the robot finishes the task (considered a success) or a maximum rollout length is reached (considered a failure). Therefore, in the generated rollouts, failed ones always have the maximum length, but successful ones are shorter. This could result in an unfair advantage for some of the compared failure detectors (if a failure detector simply learns to count the time elapsed, i.e., $s_{t}=t$, it will achieve perfect failure detection since failed rollouts have a fixed and longer duration). To ensure a fair comparison, for evaluation in Table˜1, we compute the minimum rollout length for each task and use that as $T$ for that task. The failure detection performance (in ROC-AUC) is then determined based on $s_{T}$, where $T$ is the same for all successful and failed rollouts within each task.
@@
On this benchmark, we adopt the pretrained model checkpoints of $\pi_{0}^{*}$ [64]. Note that $\pi_{0}^{*}$ model checkpoints are trained separately on the Google Robot embodiment and the WidowX embodiment, which results in two model checkpoints that have different internal feature spaces. Therefore, all failure detectors are trained and evaluated on each embodiment separately as well. All reported evaluation metrics are computed separately for each embodiment and then averaged. In Table˜2, we list the tasks used for failure detection on SimplerEnv. We generate 100 rollouts for each task with random initial configurations, and the success rates of $\pi_{0}^{*}$ on each task are also listed in Table˜2. A rollout stops after the maximum number of allowed timesteps have passed, regardless of task success or failure. Within each embodiment, 1 out of 4 tasks is unseen, and within the seen tasks, 66% of the rollouts are in $\mathcal{D}_{\text{train}}$ and the remaining 33% in $\mathcal{D}_{\text{eval-seen}}$.
Table 2: List of tasks used in $\pi_{0}^{*}$ + SimplerEnv benchmark.
@@
Real-world experiments with Franka robot: In Table˜3, we list the tasks used in the real-world experiments. For each task, we set a number of timesteps $T$ allowed for one rollout, and all rollouts of the same task are terminated after the same $T$ timesteps regardless of task success or failure. In Fig.˜7, we further visualize some example successful and failed rollouts from the real-world experiments.
Table 3: List of tasks used in the real-world Franka experiments.
@@
Real-world experiments with WidowX robot: We also tested the OpenVLA model pretrained on the “Open-X Magic Soup++" dataset on a real WidowX robot arm in our lab. In this experiment, we collected a total of 532 rollouts on the following 8 tasks (listed in Table˜4), with 244 successes and 288 failures. Each task has roughly the same number of rollouts.
Table 4: List of tasks used in the real-world experiments.
Task Instruction
1 Lift AAA Battery
2 Lift Eggplant
3 Lift Red Bottle
4 Lift Blue Cup
5 Put Blue Cup on Plate
6 Put the Red Bottle into Pot
7 Put the Carrot on Plate
8 Put the Red Block into the Pot
@@
B.6 Benchmark Statistics
Benchmark Number of Tasks Number of rollouts Seen Unseen Total Train Eval Seen Eval Unseen Total LIBERO 7 3 10 210 140 150 500 $\pi_{0}^{*}$ SimplerEnv, Google Robot 2 2 4 198 102 100 400 $\pi_{0}^{*}$ SimplerEnv, WidowX 2 2 4 198 102 100 400 Octo SimplerEnv 9 3 12 594 306 300 1200 Real Franka 10 3 13 450 150 180 780 Real WidowX 6 2 8 250 133 149 532
Table 5: Benchmark statistics for how tasks and rollouts are split into different subsets. Note that as we preformed multiple runs with different random seeds for all experiments, each run will use different set of tasks for seen and unseen subsets.
In the Table˜5, we summarize the statistics on the number of tasks and rollouts collected for each benchmark and how they are split into training and evaluation subsets. We note that as SAFE is designed for multitask failure detection, it is trained on only a limited set of training tasks and rollouts and can generalize to new tasks without further collecting rollouts. While SAFE does require hundreds of rollouts from multiple tasks during training, when handling new tasks, SAFE becomes more efficient than existing task-specific failure detectors (like [8, 18]) that require collecting rollouts for training and calibration for every new task encountered.
B.7 Training Details
@@ B.8 Hyperparameter Tuning
To determine the hyperparameters for the proposed SAFE and baselines, we perform a grid search over them and select the ones with the highest failure detection performance (ROC-AUC) on the $\mathcal{D}_{\text{eval-seen}}$ split. In Table˜9, Table˜10, and Table˜11, we report the hyperparameters we have searched over and the values with the best performance. Note that for the real-world experiments, we fix the $\mathbf{e}_{t}$ to be the “pre-logits” with “Mean” aggregation.
Appendix C Additional Results
C.1 Feature Visualization and Analysis
@@
We perform the feature analysis similar to Section˜4.1 and Fig.˜1 on other benchmarks and show the plots in Fig.˜8, Note that in this feature analysis process, the t-SNE algorithm was performed on the VLA’s embeddings without any learning. Therefore, the feature dimension reduction process is unsupervised and does not know about task successes or failures.
@@
Comparing the plots in Fig.˜1 and Fig.˜8, we can see that the embedding spaces from VLAs are different from each other, which corresponds to the different failure patterns presented by the VLAs. For $\pi_{0}$-FAST (Fig.˜1) and $\pi_{0}$ on LIBERO (Fig.˜8a and b), when task execution fails, the embeddings fall into the same region (“failure zone”). This corresponds to the major failure mode of $\pi_{0}$-FAST trained on the LIBERO dataset, where the predicted actions $\mathbf{A}_{t}$ become unstable and the robot arms move to weird configurations and out of the observation frame. For OpenVLA on LIBERO (Fig.˜8c and d). we observe that for most failed rollouts, the robot freezes at or shakes around certain configurations during the middle of task execution. Such failed rollouts result in features very close to each other, which corresponds to small blobs of red dots in Fig.˜8c.
@@
Despite the different appearances of the embedding spaces from the above benchmarks, their successful and failed rollouts are separable in the feature space. This is aligned with the high performance of the proposed SAFE and the embedding-based baseline methods. Moreover, although the embeddings of the failed rollouts from OpenVLA are spread over the space and do not form a unified “failure zone”, SAFE is still able to learn to separate task failures from successes (possibly by extracting the correlations that are not visualized by t-SNE) and generalize well to unseen tasks, as reported in Table˜1.
@@
However, the visualized embeddings of $\pi_{0}$-FAST on the real Franka robot (Fig.˜8e and f) are different, where embeddings from successful and failed rollouts are not easily separable through the t-SNE visualization. We hypothesize that because the tasks we used for real-world experiments are more diverse, their failures do not have a unified semantic meaning, and thus the embeddings are not clearly separated in the visualization. This explains the limited performance of all failure detection methods as reported by Fig.˜6, where ROC-AUC is at most 64 on $\mathcal{D}_{\text{eval-unseen}}$. Nevertheless, SAFE-MLP still outperforms all baselines on both seen and unseen splits in this evaluation.
@@
C.2 Conformal Prediction Results
@@
We use functional CP [8, 21] to determine the time-varying thresholds $\delta_{t}$ for failure detection. By varying the significance level $\alpha$ used in functional CP, we can adjust the conservativeness of failure detection and get different performance. In Fig.˜9, we plot the change of TNR (True Negative Rate), TPR (True Positive Rate) and Bal-acc (Balanced Accuracy, $\frac{\text{TNR}+\text{TPR}}{2}$) w.r.t. $\alpha$. From Fig.˜9, we can observe that while the metrics do vary with the $\alpha$, choosing $\alpha=0.15$ (or in general, between 0.05 and 0.2) performs well across the board. We have also chosen $\alpha$ to be 0.15 for most qualitative results and analyses reported in the paper.
@@
Note that we calibrate the CP bands on successful rollouts (negative data points), and thus if the assumptions used in CP ($s_{t}$ are sampled i.i.d.) hold, the TNR rate is lower bounded by and close to $1-\alpha$ (the gray dashed line in the TNR plots in Fig.˜9[21]. However, as the multitask failure detection problem requires detecting failures on tasks that are not in the training or the calibration sets, we need to calibrate CP bands on $\mathcal{D}_{\text{eval-seen}}$ and then evaluate them on $\mathcal{D}_{\text{eval-unseen}}$. Therefore, the i.i.d assumption may not hold, and TNR may deviate from the gray dashed line.
@@
From Fig.˜9, we can see that on OpenVLA+LIBERO and $\pi_{)}^{*}$+SimplerEnv benchmarks, the TNR curves obtained by SAFE are close to the gray dashed line $1-\alpha$, while those on the other 3 benchmarks are lower than $1-\alpha$. A similar phenomenon is also observed for the baseline methods: none of the TNR curves obtained from the baselines consistently conform to the $1-\alpha$ curve across all benchmarks. We attribute this to the challenging nature of the multitask failure detection problem, where the failure scores for calibration and evaluation may not come from the same distribution. Nevertheless, we still adopt the functional CP as a principled method to determine the time-varying failure detection threshold $\delta_{t}$. Moreover, from Fig.˜9, we can see that SAFE can achieve higher TPR and result in fewer false negatives compared to the baselines. This is crucial for safety-critical environments, where a missing failure (false negative) can be much more catastrophic than a false alarm (false positive).
@@
Figure 9: Additional failure detection results using $\delta_{t}$ obtained by functional CP. These plots show TNR (left column), TPR (middle column), and Bal-acc (right column) w.r.t. the significance level $\alpha$, for each evaluation benchmark. These plots are obtained with random seed=0=0.
C.3 Failure Detection Time
@@
As mentioned in Section˜6.2, we manually label when the failure happens for the failed rollouts. The labeling process is based on video recordings after all rollouts are collected and no interventions were done during the task execution. While the exact times of failure are clear for some failure modes (e.g. dangerous actions, breaking objects), they can be ambiguous and hard to annotate for other failure modes. For example, a policy may freeze in the middle of task execution, and after that either recovering from it or getting stuck indefinitely can be possible. In another case, a policy may repeatedly try grasping the object but keep missing the grasp until timeout, and it’s hard to determine a single point of failure. To handle such cases, we instruct the human annotators to pick the time where they think intervention is needed and they should take over control to prevent an execution failure. In practice, for the above ambiguous failure modes, we annotate the failures after the policy gets stuck by a few seconds or re-tries the grasping action a few times. For some rollouts that look very plausible but do not succeed due to the time limit, the failure time is annotated as the end of the rollout. Note that we annotate only the failed rollouts and not the successful ones, even though they may also show subtle signs of failure in the middle.
@@
In Fig.˜10, we compare the times of failure detected by the proposed SAFE-MLP model and a human annotator. From Fig.˜10, we can see that for both $\pi_{0}$ and $\pi_{0}$-FAST models, SAFE-MLP can detect failures before they happen (as identified by a human). When used for $\pi_{0}$-FAST deployed on LIBERO, SAFE-MLP can forecast failures well in advance and even predict 40% of the failures after the first timestep.
@@
Furthermore, from Fig.˜10a and Fig.˜10c, we can see that the blue curves jump up on the right edge of the plots. This means that the human annotator does not think these rollouts are failures until the very last moment, where the VLA model is probably on the right track and fails only due to timeout. We think such failures are also hard for failure detectors to detect, and it explains the low performance of all failure detectors on these benchmarks.
@@ C.4 Result Variance
In Table˜8, we report the standard deviation for all results in Table˜1 and Fig.˜6 left. Note that for the repeated runs, not only are they using different random seeds, also the tasks are split differently into the seen and the unseen subsets. Since different tasks have different difficulties for failure detection, it is normal to see large standard deviations in Table˜8. From Table˜8, we can see that the proposed SAFE methods achieve high averaged performance with relatively low standard deviations compared to the baselines, across all evaluation benchmarks. This signifies the strong and also stable performance of SAFE.
Appendix D Additional Ablation Studies
D.1 Number of Training Tasks
As SAFE learns to distinguish failures from successes from training rollouts, the diversity of failure modes and the number of tasks in the training data have an effect on the failure detection performance. To quantify this effect, we conduct an experiment varying the number of seen tasks that are used in training. Note that different tasks typically also have different failure modes, and in this way, we are also ablating the diversity of failure modes.
In Table˜6, we report the failure detection ROC-AUC on the OpenVLA+LIBERO benchmark, trained on different numbers of tasks. While the number of seen tasks is ablated, all experiments use the same set of unseen tasks for evaluation, and performance on unseen tasks is comparable. All numbers are averaged over 3 random seeds. Experiments with 7 tasks for training match the setting reported in the paper. Training-free methods do not depend on training tasks and are not shown.
Table˜6 shows that for most methods, with more tasks used for training, the performance on unseen tasks gets better. The proposed SAFE-MLP performs well in all settings and can also achieve good performance when fewer (3 or 5) tasks are used for training.
Table 6: Performance on the OpenVLA+LIBERO benchmark using different numbers of training tasks.
# Training Tasks 1 3 5 7 Eval Task Split Seen Unseen Seen Unseen Seen Unseen Seen Unseen Mahalanobis 40.21 52.75 58.00 52.31 57.68 50.78 62.03 58.85 Euclid. $k$-NN 49.74 63.76 61.66 67.02 59.14 67.11 66.00 55.23 Cosine. $k$-NN 53.27 60.76 65.39 65.64 67.46 70.57 67.09 69.45 PCA-KMeans 60.39 40.58 61.18 52.87 61.50 53.06 57.18 55.10 RND 29.29 50.32 54.46 47.39 56.71 49.15 52.57 46.88 LogpZO 61.75 56.17 52.89 50.49 65.99 56.60 61.57 52.91 SAFE-LSTM 50.88 52.25 68.85 63.31 70.70 66.31 70.24 72.47 SAFE-MLP 54.34 63.76 67.86 67.03 69.32 68.17 72.68 73.47
D.2 Features from Foundation Models
In Table˜7, we ablate SAFE-MLP and SAFE-LSTM using DINOv2 features, CLIP features, DINOv2 and CLIP concatenated (DINOv2+CLIP), and the VLA last-layer features (VLA; our main method). DINOv2 and CLIP features are extracted from the observation images, and this experiment is conducted over the real-world Franka rollouts, following the same setting as reported in the paper. Numbers are averaged ROC-AUC on the Seen and Unseen tasks.
The best performing method in the above table is the SAFE-MLP method based on VLA last-layer features, where VLA features outperform other feature types by a large margin. We think that this is because VLA feature space learns high-level information about the tasks, and thus can more easily distinguish failures from successes than general pretrained models. Similar findings were also reported in related works like [18].
Table 7: Comparison of model performance across different visual encoders and architectures.
Method LSTM MLP
Eval Task Split Seen Unseen Seen Unseen
DINOv2 76.93 56.96 76.20 59.46
CLIP 76.77 52.71 77.88 59.77
DINOv2+CLIP 77.09 59.65 76.36 58.43
VLA (Ours) 77.27 58.70 86.76 64.16
Appendix E Additional Discussions
E.1 Comparing Failure Detection, Uncertainty Quantification and OOD Detection
Failure detection, uncertainty quantification and OOD detection are three closely connected concepts with subtle differences. SAFE learns to model the probability of failures and detect failures of VLA policies, but it achieves this not through uncertainty quantification (UQ) or OOD detection. Here, we provide a detailed discussion comparing these three concepts.
Failure detection is the task of detecting failures when a robot is performing certain tasks. SAFE learns the likelihood of failure through training on a set of successful and failed rollouts. SAFE-LSTM is trained by BCE loss and outputs a normalized score indicating the probability of failure of VLA. The output of SAFE-MLP is not normalized and thus not a probability. However, output scores from both SAFE-LSTM and SAFE-MLP are calibrated through functional Conformal Prediction CP and can be used for failure detection with theoretical guarantees.
Uncertainty quantification (UQ) measures a VLA’s uncertainty in its outputs and can be used as a proxy for failure detection. In our experiments, the token uncertainty baselines and sample consistency baselines are inspired by LLM/VLM literature and designed based on UQ. Methods in this category are typically training-free, but they only show limited success according to our experiments.
OOD detection-based failure detection methods treat successful rollouts as normal execution conditions and assume that deviations from this norm lead to a higher chance of failure. In our experiments, the embedding distribution-based baselines are designed to detect policy failure based on OOD detection. Methods in this category can work without failed rollouts. In our experiments, we adapted them to take in both successes and failures, and they showed strong performance. Please see Section˜2.2 in our paper and also [8] for more comprehensive discussions on these methods.
Uncertainty quantification methods have been widely used for LLM/VLM hallucination detection (see Section˜2.3), and OOD detection-based methods have been shown to be effective for failure detection in robotics policies (see Section˜2.2). Therefore, we think it’s appropriate to use them as baselines.
Different from existing works based on UQ or OOD detection, SAFE directly learns to detect failures from a history of observations and the language instruction specifying the desired task without using uncertainty or OOD detection as the proxy measurement. Experiments show that this direct learning regime used by SAFE is more effective and achieves better performance than other methods.
Appendix F Potential Future Works
F.1 Extending Beyond the Last-layer features
In this paper, we maximize the simplicity and transferability when designing SAFE. By only taking feature vectors at the last layer, the proposed method can be easily integrated into any VLA models with minimal implementation changes and no finetuning on the VLAs themselves.
Further fusing or aggregating deep features from multiple layers can also benefit failure detection and is a promising future direction. Related works have shown potential in this direction. For example, [55] proposed Truthfulness Separator Vector (TSV), which is injected in the LLM latent features in the middle of the transformer and is optimized to better separate the hallucinated and truthful responses in the final token feature space. We think a similar technique can also be developed for VLA failure detection. However, this would require a special design and implementation for each VLA model (as some VLA models output discrete tokens, and others use flow matching to output continuous actions, there may not be a single design that can be applied to all VLA architectures), reducing its transferability. We leave the development of such methods as a promising future work.
Using the latent feature from an intermediate transformer block may also be a promising future direction. As shown by [49] and [53], LLM latent features from different layers have different performance on hallucination detection, and the best one may not be the last layer. However, exactly which layer works the best may depend on the model and require extensive ablation experiments to find out. For example, as reported by [49], for the OPT-6.7B model, the 20th layer works the best, but for LLAMA2-7B, the 16th layer works the best. To locate the best layer, [49] has to perform a grid search over each LLM tested. On the contrary, in our setting, VLA users can avoid such grid search experiments and simply choose the final layer for failure detection. Therefore, we think that precisely finding the layer that works the best for VLA failure detection is outside the scope of this paper, but it would be very interesting to explore for future work.
F.2 Adaptive Thresholding by Online Conformal Prediction
The proposed SAFE and baselines can be extended to online or adaptive conformal prediction (CP) [75]. In such a framework, rollout results are observed one-by-one and compared to the prediction results, and the significance level $\alpha$ is adjusted for each individual task if the prediction is wrong. However, when VLA policies are deployed, they may constantly meet novel environments and customized task instructions, and may rarely repeat the same task. In such a case, it may be less pratical to develop adaptive CP band for each specific task. Therefore, in this work, we focus on the performance of failure detectors when they are directly deployed on a novel task, and have never seen or repeated the task before. In this setting, offline CP is more appropriate. Nevertheless, we believe online CP is an interesting extension to our work, and we leave it as an important future work.
F.3 Using Detected Failures for Behavior Improvement
SAFE focuses on detecting failures accurately and in a timely manner, which enables either the robot to abort potentially dangerous actions or a human monitor to step in and take over control. How to learn a recovery policy or how to improve the VLAs themselves are important areas for future work. In this work, we focused on detecting failures of multitask VLAs reliably and in real-time, which is a crucial stepping stone towards autonomous recovery (e.g., with a fallback policy) and policy improvement (e.g., through interactive imitation learning).
We think it is possible to use the findings from this paper to further develop methods for steering or improving VLA behaviors. For example, we show that the embeddings for successful and failed rollouts are separated in the latent space, so it is possible to learn a steering vector that manipulates the latent activations of a VLA and changes its output actions, as done in [50] and [76]. However, different from the stylization or hallucination reduction tasks for LLMs, robot manipulation involves multistep closed-loop interaction between the policy and the environment, which greatly complicates the relationship between VLA outputs and task execution successes. Therefore, how to improve VLAs through activation steering is a challenging and open research question beyond the scope of our paper.